A market consists of two individuals. Their demanded equations are Q1 = 16- 4P and Q2 = 20 –2P, respectively.

What is the market demand equation?

At a price of $2, what is the point price elasticity for each person and for the market?
1
Expert's answer
2020-11-02T10:20:40-0500

Market demand equation:

The market demand is an aggregation of individual demand functions. Therefore, assuming market demand function is given by Q,

"Q = Q_{1} + Q_{2}"

"= (16-4P)+(20-2P)"

"=16+20-4P-2P"

"\bold {Q=36-6P}" "(Answer)"


Point price elasticity of demand:

When "P =\$2,"

"Q_{1}=16-4(2)"

"= 16-8"

"= 8 \space units"


"Q_{2} = 20-2(2)"

"= 20-4"

"=16" "units"


"Q=36-6(2)"

"= 36-12"

"= 24 \space units"


Differential calculus is applied to find the point price elasticity "(\eta_{p})" using the formula,

"\eta_{p}= \dfrac {dQ}{dP}× \dfrac {P}{Q}"


For individual 1

"\dfrac {dQ}{dP} = \dfrac{d}{dP}(16-4Q)"


"= -4"


"\therefore \space \eta_{p} = -4×\dfrac {2}{8}"

"=\bold {-1(unitary)}"


For individual 2

"\dfrac {dQ}{dP} = \dfrac{d}{dP}(20-2Q)"


"=-2"


"\therefore \space \eta_{p} = -2×\dfrac {2}{16}"

"= \bold {-0.25(inelastic)}"



For the market

"\dfrac {dQ}{dP} = \dfrac{d}{dP}(36-6Q)"


"=-6"



"\therefore \space \eta_{p}=-6×\dfrac{2}{24}"


"= \bold {-0.5(inelastic)}"